measure of angle 1=2x-3 and measure of angle 2 = x +42
By solving for x, find measure of angle 3

Answer by Alan3354(21581)

measure of angle 1=2x-3 and measure of angle 2 = x +42
By solving for x, find measure of angle 3
-------------------
Call them A, B and C for 1, 2 and 3
A+B+C = 180
2x-3 + x+42 + C = 180
3x + 39 + C = 180
3x + C = 141
C = 141 - 3x
---------------
That's all that can be done without additional info.
Did you leave something out?